# CoLoRA: Continuous low-rank adaptation for reduced implicit neural modeling of parameterized partial differential equations

### v1

@@ 1 Introduction
Many phenomena of interest in science and engineering depend on physics parameters $\boldsymbol{\mu}$ that influence the temporal and spatial evolution of the system such as the Reynolds number in fluid mechanics and conductivity coefficients in heat transfer. Rapidly simulating physical phenomena for a large sample $M\gg 1$ of physics parameters $\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{M}$ is paramount in science and engineering, e.g., for finding optimal designs, inverse problems, data assimilation, uncertainty quantification, and control. Numerically solving the underlying parameterized partial differential equations (PDEs) with standard numerical methods [43, 59] for large numbers of different physics parameters is prohibitively expensive in many applications and thus one often resorts to reduced modeling.
@@
Reduced models exploit structure in PDE problems to more efficiently approximate solution fields. The conventional structure that is leveraged is low rankness in the sense of classical principal component analyses. Reduced models based on such low-rank approximations can achieve exponentially fast error decays $\mathrm{e}^{-cn}$ with the rank $n$ for a wide range of (mostly nicely behaved elliptic) problems [64, 21]. However, linear low-rank approximations are affected by the so-called Kolmogorov barrier, which states that there are classes of PDEs for which linear approximations have an error decay rate of at best $1/\sqrt{n}$ [70, 38]. Examples of PDE classes that are affected by the Kolmogorov barrier are often describing transport-dominated phenomena such as strongly advecting flows and wave-like behavior; see [74] for a survey.
@@
In this work, we build on LoRA (Low Rank Adaptation) for developing parameterizations for reduced models [41]. LoRA has been developed for fine-tuning large language models and leverages the observation that fine-tuning objectives can be efficiently optimized on low-dimensional parameter spaces [60, 1, 41]. This has been observed not only for large language models but also when approximating solution fields of physical phenomena [7, 37, 10]. We build on the pre-training/fine-tuning paradigm of LoRA but modify it to Continuous LoRA (CoLoRA) that reflects our PDE setting by allowing continuous-in-time adaptation (“fine-tuning”) of parts of the low-rank components as the solution fields of the PDEs evolve; see Figure 1. This inductive bias is in agreement with how typically physical phenomena evolve over time, namely smoothly and along a latent low-dimensional structure. By composing multiple CoLoRA layers in a deep network, we obtain a nonlinear parameterization that can circumvent the Kolmogorov barrier while having a pre-training/fine-tuning decomposition with an inductive bias that reflects the special meaning of the time variable $t$ in PDE problems; see Figure 2.
@@
Time enters CoLoRA models via the low-rank network weights that are adapted in the online (“fine-tuning”) phase of reduced modeling, rather than as input to the network as the spatial coordinates. Treating time separately from inputs such as the spatial coordinates aligns well with approximating PDE solution fields sequentially in time, rather than globally over the whole time-space domain. The sequential-in-time approximation paradigm of CoLoRA allows keeping the number of offline and online parameters low, because the solution field is approximated only locally in time. Requiring a low number of online and offline parameters indicates that the inductive bias induced by separating out the time variable from the spatial coordinate is in agreement with many physics problems. Furthermore, the CoLoRA architecture can be trained on low numbers of training trajectories compared to, e.g., operator learning [61, 62], which is important because standard numerical simulations can be expensive and thus only a limited number of training trajectories are available in many cases. Having network weights that depend on time allows combining CoLoRA with equation-driven variational approaches [55, 25, 4, 15] to obtain reduced solutions that are Galerkin-optimal, which opens the door to analyses, error bounds, and goes far beyond purely data-driven forecasting.
Figure 1: LoRA fine-tunes networks to downstream tasks by adapting low-rank matrices $\boldsymbol{A},\boldsymbol{B}$. Our CoLoRA introduces a scaling $\alpha(t,\boldsymbol{\mu})$ on the low-rank matrix $\boldsymbol{A}\boldsymbol{B}$ to adapt networks continuously to predict PDE solution trajectories.
@@
There are purely data-driven surrogate modeling methods such as operator learning [61, 62, 13] that can require large amounts of data because they aim to learn a generic operator map over the full model space. Model reduction [5, 84, 9, 6, 54] considers a more structured problem via the physics parameter $\boldsymbol{\mu}$, for which nonlinear model reduction methods based on autoencoders have been presented in [56, 57, 50, 82]; however, they can require going back to the the high-fidelity numerical model to drive the dynamics, which can be expensive. Alternative approaches learn the low-dimensional latent dynamics [32, 58, 93]. Additional literature of nonlinear model reduction is reviewed in Appendix A. There is a range of methods based on implicit neural representations that updates the network parameters either based on the equation [18, 19] or via hyper-networks [71, 96]; these methods are closest to CoLoRA, except that we will show that the CoLoRA’s low-rank adaptation achieves lower errors with lower parameter counts in our examples. There are pre-training/fine-tuning techniques for global-in-time methods such as physics-informed neural networks [78] that are purely data-driven once trained. Some of these approaches use hyper-networks [22, 20] and other meta-learning such as [30]. But typically these approaches are over time-space domains and thus do not make a special treatment of time or the PDE dynamics, which is a key feature of CoLoRA. Adaptive low-rank approximations have been used in scientific computing for a long time [53, 86, 29]; however, they use one layer only and adapt the low-rank matrices directly with time. We have a more restricted adaptation in time that has fewer parameters, which is sufficient due to the non-linear composition of multiple layers. Other low-rank approximations have been widely used in the context of deep networks [85, 98, 99, 48, 87] but not in PDE settings.
@@ 2 Parameterized PDEs
Figure 2: Left: Shows that CoLoRA’s latent states $\boldsymbol{\phi}(t;\boldsymbol{\mu})$ adapts smoothly over time (RDE example). Middle: Training a CoLoRA model with a $q=2$-dimensional latent state on the Burgers’ example gives the first latent component corresponding to translation in time and the second one to the viscosity $\mu$. Right: CoLoRA learns a continuous region of low PDE residual along which the latent trajectories evolve (Vlasov example); see Appendix B.
@@
where $u_{0}$ is the initial condition and $f$ can include partial derivatives of $u$ in $\boldsymbol{x}$. In the following, we always have appropriate boundary conditions so that the PDE problem (1) is well posed. The physics parameter $\boldsymbol{\mu}$ can enter the dynamics via $f$ and the initial condition $u_{0}$. Standard numerical methods such as finite-element [43] and finite-volume [59] methods can be used to numerically solve (1) to obtain a numerical solution $u_{\text{F}}(\cdot,\cdot;\boldsymbol{\mu}):\mathcal{T}\times\Omega\to\mathbb{R}$ for a physics parameter $\boldsymbol{\mu}\in\mathcal{D}$.
Computational procedures for learning reduced models [5, 84, 9, 6, 54] are typically split into an offline and an online phase: In the offline (training) phase, the reduced model is constructed from training trajectories
@@
3 CoLoRA networks
We introduce CoLoRA deep networks that (a) provide nonlinear parameterizations that circumvent the Kolmogorov barrier of linear model reduction and (b) impose an inductive bias that treats the time variable $t$ differently from the spatial variable $\boldsymbol{x}$, which reflects that time is a special variable in physics. In particular, CoLoRA networks allow a continuous adaptation of a low number of network weights over time to capture the dynamics of solution fields (“fine-tuning”) for different physics parameters.
@@ 3.1 LoRA layers
CoLoRA networks are motivated by LoRA [41], a method that has been introduced to fine-tune large language models on discrete downstream tasks. LoRA layers are defined as
@@ 3.2 The CoLoRA layer
Models with low intrinsic dimension are very common not only in large language models but also in many applications in science and engineering with phenomena that are described by PDEs [7, 37, 10]. However, in the PDE settings, we have the special time variable $t$ that requires us to “fine-tune” continuously as the PDE solution trajectories evolve; see Figure 1. Additionally, time imposes causality, which we want to preserve in CoLoRA models.
@@
with activation function $\sigma$ and $\mathcal{C}_{i}(\boldsymbol{x})=\boldsymbol{W}_{i}\boldsymbol{x}+\alpha_{i}(t,\boldsymbol{\mu})\boldsymbol{A}_{i}\boldsymbol{B}_{i}+\boldsymbol{b}_{i}$ for $i=1,\dots,L$. The online parameters are given by the vector $\boldsymbol{\phi}(t,\boldsymbol{\mu})=[\alpha_{1}(t,\boldsymbol{\mu}),\dots,\alpha_{q}(t,\boldsymbol{\mu})]\in\mathbb{R}^{q}$ with $q=L$ in the example (5). We will later refer to $\boldsymbol{\phi}(t,\mu)$ as the latent state. All other CoLoRA parameters that are independent of time $t$ and physics parameter $\boldsymbol{\mu}$ are trainable offline and collected into the offline parameter vector $\boldsymbol{\theta}\in\mathbb{R}^{p}$ of dimension $p\gg q$.
We note in principle $\boldsymbol{A}\boldsymbol{B}$ could be full rank without increasing the size of $q$. But this would increase the number of parameters in $\boldsymbol{\theta}$. Additionally, the authors of LoRA [41] observed that full rank fine-tuning updates under-perform low rank ones despite having more degrees of freedom, which is also in agreement with the very low ranks used in dynamic low-rank and online adaptive methods in model reduction [53, 86, 29, 73, 76, 89].
A CoLoRA network defines a function $\hat{u}:\Omega\times\mathbb{R}^{p}\times\mathbb{R}^{q}\to\mathbb{R}$ that depends on an input $\boldsymbol{x}\in\Omega$, which is the spatial coordinate in our PDE setting, the offline parameters $\boldsymbol{\theta}\in\mathbb{R}^{p}$ that are independent of time $t$ and physics parameter $\boldsymbol{\mu}$ and the online parameters or latent state $\boldsymbol{\phi}(t,\boldsymbol{\mu})\in\mathbb{R}^{q}$ that depends on $t$ and $\boldsymbol{\mu}$. A CoLoRA network can also output more than one quantity by modifying the output layer, which we will use for approximating systems of PDEs in the numerical experiments. A CoLoRA network ${\hat{u}}$ is a implicit neural representation [90, 71] in the sense that the PDE solution field is given implicitly by the parameters $\boldsymbol{\theta}$ and $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ and it can be evaluated at any coordinate $\boldsymbol{x}\in\Omega$, irrespective of discretizations and resolutions used during training.
@@
Other works have examined similar weight matrix decomposition to separate out a set of adaptable parameters in the context of PDEs [20, 95] and in implicit neural representations [49]; however, none directly treat the low rank adaptation as a function of time, which is key for our approach.
@@ 3.3 CoLoRA networks can circumvent the Kolmogorov barrier
CoLoRA networks can be nonlinear in the online parameter $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ and thus can achieve faster error decays than given by (linear) Kolmogorov $n$-widths. We give one example by considering the linear advection equation $\partial_{t}u(t,x)+\mu\partial_{x}u(t,x)=0$ as in [70] with initial condition $u(0,x)=u_{0}(x)$ and solution $u(x,t;\mu)=u_{0}(x-t\mu)$, which can lead to a slow $n$-width decay of $1/\sqrt{n}$ for a linear parameterization with $n$ parameters. In contrast, with the CoLoRA MLP network ${\hat{u}}(x;\boldsymbol{\theta},\boldsymbol{\phi}(t,\mu))=c(\mathcal{C}_{1}(\sigma_{1}(\mathcal{C}_{2}(\sigma_{2}(\mathcal{C}_{3}(x))))))$ with $L=3$ layers, we can exactly represent translation and thus the solution $u(t,x;\mu)=u_{0}(x-t\mu)$ of the linear advection equation: Set $\boldsymbol{W}_{3}=[1,0]^{T},\boldsymbol{b}_{3}=[0,1]^{T},\alpha_{3}(t,\boldsymbol{\mu})=0$ and $\boldsymbol{A}_{3},\boldsymbol{B}_{3}$ arbitrary. Further $\boldsymbol{W}_{2}=[1,0],\boldsymbol{b}_{2}=[0],\alpha_{2}(t,\boldsymbol{\mu})=-t\boldsymbol{\mu}$ and $\boldsymbol{A}_{2}\boldsymbol{B}_{2}=[0,1]$ and $\boldsymbol{W}_{1}=[1],\boldsymbol{b}_{1}=[0],\alpha_{1}(t,\boldsymbol{\mu})=0$. If we use the known initial condition as activation function $\sigma_{1}=u_{0}$ and the identity as activation function $\sigma_{2}$ and set $c=1$, then we obtain $\hat{u}(x,\boldsymbol{\theta},\boldsymbol{\phi}(t,\boldsymbol{\mu}))=u_{0}(x-t\mu)$, which is the solution of the linear advection example above. Of course this is a contrived example but it shows how translation can be represented well by CoLoRA networks, which is the challenge that leads to the Kolmogorov barrier [74]. A more detailed treatment of the approximation theoretic properties of CoLoRA networks remains an open theory question that we leave for future work.

### v2

@@ Abstract
partial differential equations, reduced modeling, nonlinear model reduction, implicit neural representations, pre-training, fine-tuning
@@ 1 Introduction
Many phenomena of interest in science and engineering depend on physics parameters $\boldsymbol{\mu}$ that influence the temporal and spatial evolution of the system such as the Reynolds number in fluid mechanics and conductivity coefficients in heat transfer. Rapidly simulating physical phenomena for a large sample $M\gg 1$ of physics parameters $\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{M}$ is paramount in science and engineering, e.g., for finding optimal designs, inverse problems, data assimilation, uncertainty quantification, and control. Numerically solving the underlying parameterized partial differential equations (PDEs) with standard numerical methods (Hughes, 2012; LeVeque, 2002) for large numbers of different physics parameters is prohibitively expensive in many applications and thus one often resorts to reduced modeling.
Figure 1: LoRA fine-tunes networks to downstream tasks by adapting low-rank matrices $\boldsymbol{A}\boldsymbol{B}$. Our CoLoRA introduces a scaling $\alpha(t,\boldsymbol{\mu})$ on the low-rank matrix $\boldsymbol{A}\boldsymbol{B}$ to adapt networks continuously to predict PDE solution trajectories.
@@
Reduced models exploit structure in PDE problems to more efficiently approximate solution fields. The conventional structure that is leveraged is low rankness in the sense of the classical principal component analysis. Reduced models based on such low-rank approximations can achieve exponentially fast error decays $\mathrm{e}^{-cn}$ with the rank $n$ for a wide range of (mostly nicely behaved elliptic) problems (Maday et al., 2002; Cohen & DeVore, 2016). However, linear low-rank approximations are affected by the so-called Kolmogorov barrier, which states that there are classes of PDEs for which linear approximations have an error decay rate of at best $1/\sqrt{n}$ (Ohlberger & Rave, 2016; Greif & Urban, 2019). Examples of PDE classes that are affected by the Kolmogorov barrier are often describing transport-dominated phenomena such as strongly advecting flows and wave-like behavior; see (Peherstorfer, 2022) for a survey.
@@
In this work, we build on LoRA (Low Rank Adaptation) for developing parameterizations for reduced models (Hu et al., 2022). LoRA has been developed for fine-tuning large language models and leverages the observation that fine-tuning objectives can be efficiently optimized on low-dimensional parameter spaces (Li et al., 2018; Aghajanyan et al., 2021; Hu et al., 2022). This has been observed not only for large language models but also when approximating solution fields of physical phenomena (Bachmayr et al., 2017; Grasedyck et al., 2013; Berman & Peherstorfer, 2023). We build on the pre-training/fine-tuning paradigm of LoRA but modify it to Continuous LoRA (CoLoRA) that reflects our PDE setting by allowing continuous-in-time adaptation (“fine-tuning”) of parts of the low-rank components as the solution fields of the PDEs evolve; see Figure 1. This inductive bias is in agreement with how typically physical phenomena evolve over time, namely smoothly and along a latent low-dimensional structure. By composing multiple CoLoRA layers in a deep network, we obtain a nonlinear parameterization that can circumvent the Kolmogorov barrier while having a pre-training/fine-tuning decomposition with an inductive bias that reflects the special meaning of the time variable $t$ in PDE problems; see Figure 2.
@@
Time enters CoLoRA models via the low-rank network weights that are adapted in the online (“fine-tuning”) phase of reduced modeling, rather than as input to the network as the spatial coordinates. Treating time separately from inputs such as the spatial coordinates aligns well with approximating PDE solution fields sequentially in time, rather than globally over the whole time-space domain. The sequential-in-time approximation paradigm of CoLoRA allows keeping the number of offline and online parameters low, because the solution field is approximated only locally in time. Requiring a low number of online and offline parameters indicates that the inductive bias induced by separating out the time variable from the spatial coordinate is in agreement with many physics problems. Furthermore, the CoLoRA architecture can be trained on low numbers of training trajectories compared to, e.g., operator learning (Li et al., 2021; Lu et al., 2021), which is important because standard numerical simulations can be expensive and thus only a limited number of training trajectories are available in many cases. Having network weights that depend on time allows combining CoLoRA with equation-driven variational approaches (Lasser & Lubich, 2020; Du & Zaki, 2021; Anderson & Farazmand, 2022; Bruna et al., 2024; Wen et al., 2024) to obtain reduced solutions that are Galerkin-optimal, which opens the door to analyses, error bounds, and goes far beyond purely data-driven forecasting.
@@
There are purely data-driven surrogate modeling methods such as operator learning (Li et al., 2021; Lu et al., 2021; Boullé & Townsend, 2023) that can require large amounts of data because they aim to learn a generic operator map over the full model space. Model reduction (Antoulas, 2005; Rozza et al., 2008; Benner et al., 2015; Antoulas et al., 2021; Kramer et al., 2024) considers a more structured problem via the physics parameter $\boldsymbol{\mu}$, for which nonlinear model reduction methods based on autoencoders have been presented in (Lee & Carlberg, 2020, 2021; Kim et al., 2022; Romor et al., 2023); however, they can require going back to the high-fidelity numerical model to drive the dynamics, which can be expensive. Alternative approaches learn the low-dimensional latent dynamics (Fulton et al., 2019; Lee & Parish, 2021; Wan et al., 2023). Additional literature of nonlinear model reduction is reviewed in Appendix A.
There is a range of methods based on implicit neural representations that updates the network parameters either based on the equation (Chen et al., 2023a, b) or via hyper-networks (Pan et al., 2023; Yin et al., 2023); these methods are closest to CoLoRA, except that we will show that CoLoRA’s low-rank adaptation achieves lower errors with lower parameter counts in our examples. There are pre-training/fine-tuning techniques for global-in-time methods such as physics-informed neural networks (Raissi et al., 2019) that are purely data-driven once trained. Some of these approaches use hyper-networks (de Avila Belbute-Peres et al., 2021; Cho et al., 2023) and other meta-learning such as Finn et al. (2017). But typically these approaches are over time-space domains and thus do not make a special treatment of time or the PDE dynamics, which is a key feature of CoLoRA. Adaptive low-rank approximations have been used in scientific computing for a long time (Koch & Lubich, 2007; Sapsis & Lermusiaux, 2009; Peherstorfer & Willcox, 2015; Einkemmer & Lubich, 2019); however, they use one layer only and adapt the low-rank matrices directly with time. We have a more restricted adaptation in time that has fewer parameters, which is sufficient due to the nonlinear composition of multiple layers. Other low-rank approximations have been widely used in the context of deep networks (Sainath et al., 2013; Zhang et al., 2014; Zhao et al., 2016; Khodak et al., 2021; Schotthöfer et al., 2022) but not in PDE settings.
@@ 2 Parameterized PDEs
[…]
Figure 2: Left: Shows that CoLoRA’s latent states $\boldsymbol{\phi}(t;\boldsymbol{\mu})$ adapt smoothly over time (RDE example). Middle: Training a CoLoRA model with a $q=2$-dimensional latent state on the Burgers’ example gives the first latent component corresponding to translation in time and the second one to the viscosity $\mu$. Right: CoLoRA learns a continuous region of low PDE residual along which the latent trajectories evolve (Vlasov example); see Appendix B.
@@
where $u_{0}$ is the initial condition and $f$ can include partial derivatives of $u$ in $\boldsymbol{x}$. In the following, we always have appropriate boundary conditions so that the PDE problem (1) is well posed. The physics parameter $\boldsymbol{\mu}$ can enter the dynamics via $f$ and the initial condition $u_{0}$. Standard numerical methods such as finite-element (Hughes, 2012) and finite-volume (LeVeque, 2002) methods can be used to numerically solve (1) to obtain a numerical solution $u_{\text{F}}(\cdot,\cdot;\boldsymbol{\mu}):\mathcal{T}\times\Omega\to\mathbb{R}$ for a physics parameter $\boldsymbol{\mu}\in\mathcal{D}$.
Computational procedures for learning reduced models (Antoulas, 2005; Rozza et al., 2008; Benner et al., 2015; Antoulas et al., 2021; Peherstorfer, 2022; Kramer et al., 2024) are typically split into an offline and an online phase: In the offline (training) phase, the reduced model is constructed from training trajectories
@@
3 CoLoRA neural networks
We introduce CoLoRA networks that (a) provide nonlinear parameterizations that circumvent the Kolmogorov barrier of linear model reduction and (b) impose an inductive bias that treats the time variable $t$ differently from the spatial variable $\boldsymbol{x}$, which reflects that time is a special variable in physics. In particular, CoLoRA networks allow a continuous adaptation of a low number of network weights over time to capture the dynamics of solution fields (“fine-tuning”) for different physics parameters.
@@ 3.1 LoRA layers
CoLoRA networks are motivated by LoRA (Hu et al., 2022), a method that has been introduced to fine-tune large language models on discrete downstream tasks. LoRA layers are defined as
@@ 3.2 The CoLoRA layer
Models with low intrinsic dimension are very common not only in large language models but also in many applications in science and engineering with phenomena that are described by PDEs (Bachmayr et al., 2017; Grasedyck et al., 2013; Berman & Peherstorfer, 2023). However, in the PDE settings, we have the special time variable $t$ that requires us to “fine-tune” continuously as the PDE solution trajectories evolve; see Figure 1. Additionally, time imposes causality, which we want to preserve in CoLoRA models.
@@
with activation function $\sigma$ and $\mathcal{C}_{i}(\boldsymbol{x})=\boldsymbol{W}_{i}\boldsymbol{x}+\alpha_{i}(t,\boldsymbol{\mu})\boldsymbol{A}_{i}\boldsymbol{B}_{i}\boldsymbol{x}+\boldsymbol{b}_{i}$ for $i=1,\dots,L$. The online parameters are given by the vector $\boldsymbol{\phi}(t,\boldsymbol{\mu})=[\alpha_{1}(t,\boldsymbol{\mu}),\dots,\alpha_{q}(t,\boldsymbol{\mu})]\in\mathbb{R}^{q}$ with $q=L$ in the example (5). We will later refer to $\boldsymbol{\phi}(t,\mu)$ as the latent state. All other CoLoRA parameters that are independent of time $t$ and physics parameter $\boldsymbol{\mu}$ are trainable offline and collected into the offline parameter vector $\boldsymbol{\theta}\in\mathbb{R}^{p}$ of dimension $p\gg q$.
We note that in principle $\boldsymbol{A}\boldsymbol{B}$ could be full rank without increasing the size of $q$. But this would increase the number of parameters in $\boldsymbol{\theta}$. Additionally, the authors of LoRA (Hu et al., 2022) observed that full rank fine-tuning updates under-perform low rank ones despite having more degrees of freedom, which is also in agreement with the low ranks used in dynamic low-rank and online adaptive methods in model reduction (Koch & Lubich, 2007; Sapsis & Lermusiaux, 2009; Peherstorfer & Willcox, 2015; Einkemmer & Lubich, 2019; Peherstorfer, 2020; Uy et al., 2022; Singh et al., 2023).
A CoLoRA network defines a function $\hat{u}:\Omega\times\mathbb{R}^{p}\times\mathbb{R}^{q}\to\mathbb{R}$ that depends on an input $\boldsymbol{x}\in\Omega$, which is the spatial coordinate in our PDE setting, the offline parameters $\boldsymbol{\theta}\in\mathbb{R}^{p}$ that are independent of time $t$ and physics parameter $\boldsymbol{\mu}$ and the online parameters or latent state $\boldsymbol{\phi}(t,\boldsymbol{\mu})\in\mathbb{R}^{q}$ that depends on $t$ and $\boldsymbol{\mu}$. A CoLoRA network can also output more than one quantity by modifying the output layer, which we will use for approximating systems of PDEs in the numerical experiments. A CoLoRA network ${\hat{u}}$ is an implicit neural representation (Sitzmann et al., 2020; Pan et al., 2023) in the sense that the PDE solution field is given implicitly by the parameters $\boldsymbol{\theta}$ and $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ and it can be evaluated at any coordinate $\boldsymbol{x}\in\Omega$, irrespective of discretizations and resolutions used during training.
@@
Other works have examined similar weight matrix decomposition to separate out a set of adaptable parameters in the context of PDEs (Cho et al., 2023; Wen et al., 2023) and in implicit neural representations (Kim et al., 2023); however, none directly treat the low rank adaptation as a function of time, which is key for our approach.
@@ 3.3 CoLoRA networks can circumvent the Kolmogorov barrier
CoLoRA networks can be nonlinear in the online parameter $\boldsymbol{\phi}(t,\boldsymbol{\mu})$ and thus achieve faster error decays than given by (linear) Kolmogorov $n$-widths. We give one example by considering the linear advection equation $\partial_{t}u(t,x)+\mu\partial_{x}u(t,x)=0$ as in Ohlberger & Rave (2016) with initial condition $u(0,x)=u_{0}(x)$ and solution $u(x,t;\mu)=u_{0}(x-t\mu)$, which can lead to a slow $n$-width decay of $1/\sqrt{n}$ for a linear parameterization with $n$ parameters. In contrast, with the CoLoRA MLP network ${\hat{u}}(x;\boldsymbol{\theta},\boldsymbol{\phi}(t,\mu))=c(\mathcal{C}_{1}(\sigma_{1}(\mathcal{C}_{2}(\sigma_{2}(\mathcal{C}_{3}(x))))))$ with $L=3$ layers, we can exactly represent translation and thus the solution $u(t,x;\mu)=u_{0}(x-t\mu)$ of the linear advection equation: Set $\boldsymbol{W}_{3}=[1,0]^{T},\boldsymbol{b}_{3}=[0,1]^{T},\alpha_{3}(t,\boldsymbol{\mu})=0$ and $\boldsymbol{A}_{3},\boldsymbol{B}_{3}$ arbitrary. Further $\boldsymbol{W}_{2}=[1,0],\boldsymbol{b}_{2}=[0],\alpha_{2}(t,\boldsymbol{\mu})=-t\boldsymbol{\mu}$ and $\boldsymbol{A}_{2}\boldsymbol{B}_{2}=[0,1]$ and $\boldsymbol{W}_{1}=[1],\boldsymbol{b}_{1}=[0],\alpha_{1}(t,\boldsymbol{\mu})=0$. If we use the known initial condition as activation function $\sigma_{1}=u_{0}$ and the identity as activation function $\sigma_{2}$ and set $c=1$, then we obtain $\hat{u}(x,\boldsymbol{\theta},\boldsymbol{\phi}(t,\boldsymbol{\mu}))=u_{0}(x-t\mu)$, which is the solution of the linear advection example above. Note that using the initial condition as an activation function is proper in this context because the initial condition is typically given in closed form or at least can be evaluated over $x$ and thus can be fitted during the pre-training.
Of course this example with the linear advection equation is contrived but it shows how translation can be represented well by CoLoRA networks, which is the challenge that leads to the Kolmogorov barrier (Peherstorfer, 2022). A more detailed treatment of the approximation theoretic properties of CoLoRA networks remains an open theory question that we leave for future work.